A population grows from 11000 to 15000 in five years. Assuming the growth is exponential, find the:
(a) Annual growth rate
(b) Continuous growth rate
Round your answers to 3 decimal places.
(a) The annual growth rate is %.
(b) The continuous growth rate is %.

Answer by stanbon(48535)

A population grows from 11000 to 15000 in five years. Assuming the growth is exponential, find the:
---------------
Exponential Form: y = ab^x
15000 = 11000b^5
b^5 = 15/11
b = 1.064
-----------------
EQUATION: y = 11000*1.064^t
----------------------------------
(a) Annual growth rate: 0.064 or 6.4%
(b) Continuous growth rate: not sure what you mean by this.
=============================================================
Cheers,
Stan H.